# Interactive Public Transport Infrastructure Analysis through Mobility Profiles: Making the Mobility Transition Transparent

### v1

@@ Abstract
Efficient public transport systems are crucial for sustainable urban development as cities face increasing mobility demands. Yet, many public transport networks struggle to meet diverse user needs due to historical development, urban constraints, and financial limitations. Traditionally, planning of transport network structure is often based on limited surveys, expert opinions, or partial usage statistics. This provides an incomplete basis for decision-making. We introduce an data-driven approach to public transport planning and optimization, calculating detailed accessibility measures at the individual housing level. Our visual analytics workflow combines population-group-based simulations with dynamic infrastructure analysis, utilizing a scenario-based model to simulate daily travel patterns of varied demographic groups, including schoolchildren, students, workers, and pensioners. These population groups, each with unique mobility requirements and routines, interact with the transport system under different scenarios traveling to and from Points of Interest (POI), assessed through travel time calculations. Results are visualized through heatmaps, density maps, and network overlays, as well as detailed statistics. Our system allows us to analyze both the underlying data and simulation results on multiple levels of granularity, delivering both broad insights and granular details. Case studies with the city of Konstanz, Germany reveal key areas where public transport does not meet specific needs, confirmed through a formative user study. Due to the high cost of changing legacy networks, our analysis facilitates the identification of strategic enhancements, such as optimized schedules or rerouting, and few targeted stop relocations, highlighting consequential variations in accessibility to pinpointing critical service gaps. Our research advances urban transport analytics by providing policymakers and citizens with a system that delivers both broad insights with granular detail into public transport services for a data-driven quality assessment at housing-level detail.
@@ Introduction
However, public transport networks often struggle to meet these increasing demands due to constraints imposed by historical development, urban infrastructure, and financial limitations [21]. Traditional planning methods for public transport systems typically rely on data sources that are inherently limited—such as small-scale surveys, expert opinions, and partial usage statistics. While complex network flow simulations exist [38, 32, 19], these primarily focus on optimizing traffic flow for a modeled infrastructure and less on typical usage patterns. Simultaneously, transport networks are frequently compromised by constraints regarding spatial limitations of dense environments, cost-benefit considerations, and persistent budgetary constraints, making this a complex optimization problem with a huge search space difficult to solve algorithmically. Due to these difficulties, these methods–while potentially providing some valuable insights–far too often result in sub-optimal planning and an incomplete understanding of network efficiencies and user needs [12].
@@ 1.1 Urban Transport Visualizations
The creation of public transport infrastructure maps is arguably one of the earliest and most prominent use cases of data visualization, as showcased by abstract representations of subway, train, or bus networks [3]. Much of the academic research on Urban Transit Routing Planning (UTRP) is focused on algorithmic routing and conducted [2] with either the openly available Multi-Agent Transport Simulation (MATSim) [38], the commercial Multimodal Traffic Simulation Software (VISSIM) [32] by PTV Group or the open-source Simulation of Urban MObility (SUMO) [19] by the German Aerospace Agency (DLR). These systems primarily focus on the algorithmic routing problem to design networks in their entirety but usually include or integrate with a basic map viewer application to visualize the network structure. Based on this, there exist multiple commercial applications to analyze public transport infrastructure and passenger data [29, 17, 18, 35, 2], regarding travel times, incidents, and real-time status of a transport system: Zeng et al. present an integrated system to visualize and explore passenger mobility data and is aimed at transportation researchers [41]. They combine an isochrone map for spatial information with an isotime map flow map for temporal information and a paired journey detail view. In our work, we go beyond the passive analysis of existing data towards generating novel data via simulation and analyzing changes to the infrastructure interactively. Burch et al. present PasVis, an interactive application to extend common schematic public transport maps with additional information in the forms of glyphs and extra views showing different statistics such as ridership at certain stations on stops [7]. We mostly utilize a layer map visualization without extra views and go beyond analyzing existing transport. Related to our objective of planning and travel time analysis, work by Baltzer et al. [5] visualizes the effect of new infrastructure on travel time. In particular, they focus on differences in travel times, represented by isochrone-based maps. We use related approaches to visualize travel times but move from static pre-computed maps to dynamic and interactive visualizations.
@@
1.2 Agent-based Simulation
@@
Similarily, Manser et al. (2020) extended an agent-based microsimulation framework to design and evaluate large-scale public transport networks for the city of Zurich, Switzerland. Their study, incorporating dynamic demand responses to network changes, suggests that a redesigned network with smaller vehicles and higher frequencies can increase transit ridership while reducing subsidies. This method also aids in identifying corridors for capacity upgrades, making it a robust tool for planning future transport infrastructure [22].
Yoon et al. (2022) developed an open-source simulation sandbox to assess various transit system designs, from fixed-route services to on-demand microtransit. By comparing these designs under different demand scenarios, their study highlights the flexibility and adaptability of semi-flexible and on-demand services, providing valuable insights for optimizing urban transit systems [40].
@@ 2.1 Underlying Data and Input
We utilize a series of public open-data sources for our analysis:
@@
The basis for route planning is formed by country-wide target timetable data. Such timetable data is usually available from local transport operators, and public authorities in a region or country often publish a compiled version in various formats, e.g., the General Transit Feed Specification (GTFS) [24]. For Germany, we use a current GTFS-NeTEx dataset [10] from the DELFI integration platform (DIP) available via an open API, providing network, timetables, routes, and stop information. It contains c. 645,000 stops and 1,600,000 daily trips from 434 transport agencies for various transport modes, including bus, subway, tram, regional and high-speed rail, and ferry, among others.
For geographical information, we leverage OpenStreetMap (OSM) dumps (*.osm.pbf) available from Geofabrik [14], providing streets, map tiles for the visualization, and VGI-POI information. While in theory, housing location data could be extracted from OSM, due to quality issues, we consider dedicated datasets like HK-DE [39] for housing coordinates. General population data for individual counties (detailing age distribution and population density), to enable demographic weighting, is retrieved from census data [36]. Figure 4 shows the map data layers that can be used to inspect the underlying source data.
@@ 2.2 Profile-Based Computation
Population groups and their associated profiles can be created arbitrarily by the users, for example, groups such as schoolchildren, students, working people, and the elderly. These mobility profiles specify daily schedules within a prototypical week and frequented points of interest for these different person groups. For the default profiles, we use statistical data obtained from large population surveys like Mobility in Germany [27], which details daily activity patterns for such groups, but this is adaptable when detailed statistics or estimates are available.
@@ 3 Visual Analytics Application
In this chapter, we present the design of a visual analytics application that enables interaction between the human, data, and simulation, enabling multiple analysis modes. In the previous chapter, we have described the underlying simulation system in the previous section, which informs three specific requirements for our agent-based public transport analytics system: (R1) Visualizing the underlying data and geospatial information informing the travel time simulation, (R2) the automated generation of a simulation based on publicly available data, and (R3) the interactive adaptation of the simulation, e.g., by placing new stops and modifying existing locations. We start with an overview of the main user interface before addressing how our design meets these requirements.
@@ 4 Case Study
Our system aims primarily at policymakers and non-specialists who have an interest in the analysis and improvement of public transport infrastructure but do not have access to more detailed closed data and do not have the resources to work with numerical flow simulation planning tools. Such users might include political activists, lobbying groups, journalists, or interested citizen scientists. We argue that a broader representation of interested and supporting tools can strengthen the quality of public discourse and decision-making.
@@ 4.2 Case 2: Evaluating the Effect of Infrastructure Changes
(G2, G3) But how could such change look like, and how can one use our approach towards modeling and recommending potential changes to the transport infrastructure? For this case study, we chose the smaller city of Konstanz, Germany, with a population of around 85.000 inhabitants.
Alex is a student at the local university and is often frustrated that his home takes a long time to reach. He wonders if there are possible improvement opportunities in his neighborhood. Similar to the previous case, Alex performs a heatmap analysis, where he notices that bus lines stop on the very edge side of his neighborhood and hypothesizes that rerouting a bus route through its center might improve travel times for many houses. After identifying possible locations, he switches to the editing mode tab (Subsection 3.4). In the editing mode, the user can choose to place elements such as POIs, bike sharing, or new transport infrastructure (Figure 6). Alex wants to evaluate the effect of a changed bus stop placed at a strategic location. He first selects the bus stop icon in the menu bar. The bus line overlay gets activated, and he can click on an existing bus line to move one stop from this bus line. After clicking on a bus line, he moves the stop from the edge to the center (Figure 6.2/3). Now, the simulation is re-computed in two stages, indicated by a loading icon: Within a region constrained by the neighboring bus stop locations, walking times to the closest location are re-computed to consider the change. Simultaneously, the travel times from the new bus stop to the relevant POIs are recomputed (i.e., a new row to the travel-time matrix is computed), as described above. Finally, for all relevant points, the travel time estimates are updated and are displayed on the travel time map (Figure 6.5). To evaluate the effect of the change, the user can select a difference visualization, i.e., a heatmap that displays the differences in travel time estimates with and without the modification (compare to Figure 7). Alex tries out multiple placements in order to identify the best possible location.
@@
4.3 Formative User Study
We conducted a formative user study with eleven participants recruited from a university environment, aged between 21 to 32 years: six computer science (CS) PhD students (PHD1-6), three CS master students (CS1-3), one finished CS master (CS4), and one history student (HS1). Of those, ten had prior knowledge of visual analytics, and seven had previously worked with GIS tools. The study was conducted in a university lab in a controlled environment, with a short preliminary questionnaire, a short introduction, a task-based assessment using think-aloud, followed by NASA TLX and User Experience Questionnaire (UEQ) as well as a semi-structured interview at the end, taking in total about one hour. The setup was shown on a 14.2-inch display with a resolution of 3024x1964, using the web-based prototype.
@@ 5 Discussion and Future Work
The case studies have shown the comparative ease with which a scenario-based analysis of existing public transport infrastructure can be generated, the effects of change can be analyzed, and the insights that can be drawn from it, which is validated by the formative user study. In the following, we want to highlight some key findings before discussing the limitations of our approach.
@@ 5.1 Findings and Lessons Learned
We outline how our system matches the three goals of (G1) visualizing the quality of existing infrastructure, (G2) enabling detailed analysis for different demographic groups and usage scenarios, (G3) simulate the approximate effectiveness of modifications in the case studies (see Section 4 and formative user study (see Section 4.3). This is achieved via simple yet expressive visualizations (R1), a simulation based on open-source real-world data (R2), which is easily adaptable (R3).
@@ 5.2 Limitations and Future Work
The modifications of the transport network does not consider secondary effects in our existing simulation framework. Right now, for example, when adding a new stop or changing the schedule, it would make sense to perform changes to timetables also of other lines, adapting their schedules, too, to ensure smooth connections and reduce waiting times. This is not fully supported, and we do not aim to perform such progressive schedule updates. This is because such changes can quickly explode and would require complex adaptation of timetables and routes, requiring an almost complete (and slow) re-simulation of the whole area, which is not compatible with the interactive approach so far. Our system could be extended to integrate other agent-based simulation frameworks like MatSim [38] or Sumo [19], in the future, one could also utilize more computationally extensive agent-based simulations to better support this.
@@
Our system is designed to be accessible and user-friendly, especially for non-expert users. While we expect some time required to become familiar with all functionalities, we have already evaluated the usability for non-expert users in a formative user study with eleven participants, indicating its usability for individuals or interest groups. The usability of the approach could be investigated in a quantitative user study. This could open up public transport planning to wider audiences and improve the quality of discussion for integrating civil society into planning processes. We can imagine our approach to enable even broader participation of citizens, e.g., we could use the mobility profile editor to collect survey data from citizens and aggregate the results to get more accurate profiles in areas where surveys have been limited so far.
The interactive nature requires some computing capability for a user, which is achievable for a domain expert setting but becomes a scalability issue when making the system open to a general audience. In the future, one could imagine a public read-only version of the system, while interactivity would be limited to a restricted user group, for example, through prior registration and application.
Due to similar reasons, multi-modal and cross-modal transport is not yet considered, for example, combining public transport with individual mobility like a car or a bicycle. Our approach so far only considers public transport in itself, which might not always reflect actual and desirable usage (e.g., taking the bike to the train station), but this could be added in the future with more compute available.
When used by city planners and other expert users, integration with de-facto standard frameworks like ArcGIS might also be desirable.

### v2

@@ Abstract
Efficient public transport systems are crucial for sustainable urban development as cities face increasing mobility demands. Yet, many public transport networks struggle to meet diverse user needs due to historical development, urban constraints, and financial limitations. Traditionally, planning of transport network structure is often based on limited surveys, expert opinions, or partial usage statistics. This provides an incomplete basis for decision-making. We introduce an data-driven approach to public transport planning and optimization, calculating detailed accessibility measures at the individual housing level. Our visual analytics workflow combines population-group-based simulations with dynamic infrastructure analysis, utilizing a scenario-based model to simulate daily travel patterns of varied demographic groups, including schoolchildren, students, workers, and pensioners. These population groups, each with unique mobility requirements and routines, interact with the transport system under different scenarios traveling to and from Points of Interest (POI), assessed through travel time calculations. Results are visualized through heatmaps, density maps, and network overlays, as well as detailed statistics. Our system allows us to analyze both the underlying data and simulation results on multiple levels of granularity, delivering both broad insights and granular details. Case studies with the city of Konstanz, Germany reveal key areas where public transport does not meet specific needs, confirmed through an initial user study. Due to the high cost of changing legacy networks, our analysis facilitates the identification of strategic enhancements, such as optimized schedules or rerouting, and few targeted stop relocations, highlighting consequential variations in accessibility to pinpointing critical service gaps. Our research advances urban transport analytics by providing policymakers and citizens with a system that delivers both broad insights with granular detail into public transport services for a data-driven quality assessment at housing-level detail.
@@ Introduction
However, public transport networks often struggle to meet these increasing demands due to constraints imposed by historical development, urban infrastructure, and financial limitations [23]. Traditional planning methods for public transport systems typically rely on inherently limited data sources—such as small-scale surveys, expert opinions, and partial usage statistics. While complex network flow simulations exist [40, 34, 21], traditional urban planning, when based on movement patterns, primarily focuses on optimizing traffic flow for a modeled infrastructure and less on typical usage patterns. Simultaneously, transport networks are frequently compromised by constraints regarding spatial limitations of dense environments, cost-benefit considerations, and persistent budgetary constraints, making this a complex optimization problem with a vast search space difficult to solve algorithmically. Due to these difficulties, these methods–while potentially providing some valuable insights–often result in sub-optimal planning and an incomplete understanding of network efficiencies and user needs [13].
@@
The user interface of our approach, enabling the exploration and interactive simulation of Mobility Profiles:
It consists of two main visual elements: a map view with a legend (a) and a sidebar (b-f) with several control components.
The background information (b), as well as particular network and POI information, can flexibly be toggled.
The scenarios reflecting the mobility of a set of demographics can be modified (d), compared to different assumptions (c), and interactively partly selected (e). Further general settings are available (f).
@@ 1.1 Urban Transport Visualizations
The creation of public transport infrastructure maps is arguably one of the earliest and most prominent use cases of data visualization, as showcased by abstract representations of subway, train, or bus networks [3]. Much of the academic research on Urban Transit Routing Planning (UTRP) is focused on algorithmic routing and conducted [2] with either the openly available Multi-Agent Transport Simulation (MATSim) [40], the commercial Multimodal Traffic Simulation Software (VISSIM) [34] by PTV Group or the open-source Simulation of Urban MObility (SUMO) [21] by the German Aerospace Agency (DLR). These systems primarily focus on the algorithmic routing problem to design networks in their entirety but usually include or integrate with a basic map viewer application to visualize the network structure. Based on this, there exist multiple commercial applications to analyze public transport infrastructure and passenger data [31, 18, 19, 37, 2], regarding travel times, incidents, and real-time status of a transport system: They combine an isochrone map for spatial information with an isotime map flow map for temporal information and a paired journey detail view. In our work, we go beyond the passive analysis of existing data towards generating novel data via simulation and analyzing changes to the infrastructure interactively. Burch et al. present PasVis, an interactive application to extend common schematic public transport maps with additional information in the forms of glyphs and extra views showing different statistics such as ridership at certain stations on stops [7]. Zeng et al. use a glyph-based visualization to uncover the relationship between human mobility and point of interests [43]. We mostly utilize a layer map visualization without extra views and go beyond analyzing existing transport. Related to our objective of planning and travel time analysis, work by Baltzer et al. [5] visualizes the effect of new infrastructure on travel time. In particular, they focus on differences in travel times, represented by isochrone-based maps. We use related approaches to visualize travel times but move from static pre-computed maps to dynamic and interactive visualizations.
@@
1.2 Public Transport Simulation
@@
Similarily, Manser et al. (2020) extended an agent-based microsimulation framework to design and evaluate large-scale public transport networks for the city of Zurich, Switzerland. Their study, incorporating dynamic demand responses to network changes, suggests that a redesigned network with smaller vehicles and higher frequencies can increase transit ridership while reducing subsidies. This method also aids in identifying corridors for capacity upgrades [24].
Yoon et al. (2022) developed an open-source simulation sandbox to assess various transit system designs, from fixed-route services to on-demand microtransit. By comparing these designs under different demand scenarios, their study highlights the flexibility and adaptability of semi-flexible and on-demand services, enabling for optimizing urban transit systems [42].
@@
Some work has used mobility profiles and points of interest data in public transport contexts, mainly to enhance existing lines and schedules, but often failing to address underserved neighborhoods or plan new lines [12]. Li et al. (2018) utilized point of interest data to identify employment locations and commuting patterns, employing a categorization similar to ours, but without directly linking it to public transport planning [20].
@@ 2.1 Underlying Data and Input
Public transport analysis can leverage a magnitude of both public and private data and it in general requires transport network and schedule information, infrastructure maps (e.g., streets), as well as flow information. We utilize a series of well-known public open-data sources for our analysis, relying on GTFS, OSM, as well infering flow based on typical interest and population densities:
@@
The basis for route planning is formed by country-wide target timetable data. Such timetable data is usually available from local transport operators, and public authorities in a region or country often publish a compiled version in various formats, e.g., the General Transit Feed Specification (GTFS) [26]. For Germany, we use a current GTFS-NeTEx dataset [10] from the DELFI integration platform (DIP) available via an open API, providing network, timetables, routes, and stop information. It contains c. 645,000 stops and 1,600,000 daily trips from 434 transport agencies for various transport modes, including bus, subway, tram, regional and high-speed rail, and ferry.
For geographical information, we leverage OpenStreetMap (OSM) dumps (*.osm.pbf) available from Geofabrik [15], providing streets, map tiles for the visualization, and VGI-POI information. While in theory, housing location data could be extracted from OSM, due to quality issues, we consider dedicated datasets like HK-DE [41] for housing coordinates. General population data for individual counties (detailing age distribution and population density), to enable demographic weighting, is retrieved from census data [38]. Figure 4 shows the map layers that can be used to inspect the underlying source data.
@@ 2.2 Profile-Based Computation
The Scenario Editor to modify the mobility profiles of different demographic groups (B).
For each demographic, different categories of interest (A, i.e., one or more different POI types) are modeled with a probability of visit (C) for a particular time of day within a prototypical week. The inset (D) shows a more detailed example.
@@
Population groups and their associated profiles can be created arbitrarily by the users, for example, groups such as schoolchildren, students, working people, and the elderly. These mobility profiles specify daily schedules within a prototypical week and frequented points of interest for these different person groups.
A more detailed Mobility Profile card example for a group Elderly
@@
For the default profiles, we use statistical data obtained from large population surveys like Mobility in Germany [29], which details daily activity patterns for such groups, but this is adaptable when detailed statistics or estimates are available.
@@
Data Exploration Mode: Map layers to inspect underlying information like transport lines, stop locations, points of interest, and housing data.
@@ 3 Visual Analytics Application
In this section, we present the design of a visual analytics application that enables interaction between the human, data, and simulation, enabling multiple analysis modes. In the previous section, we have described the underlying simulation system, which informs three specific requirements for our agent-based public transport analytics system: (R1) Visualizing the underlying data and geospatial information informing the travel time simulation, (R2) the automated generation of a simulation based on publicly available data, and (R3) the interactive adaptation of the simulation, e.g., by placing new stops and modifying existing locations. We start with an overview of the main user interface before addressing how our design meets these requirements.
The application allows a comparison of travel times for different demographics, based on generated mobility profiles. We can identify neighborhoods that are especially suited or unsuited for different demographic groups. This indicated the potential for targeted modifications.
@@ 3.4 Editing Mode
Our system enables users to interactively place transport infrastructure, POIs, and housing to explore the effects of new or updated infrastructure. In this scenario, a new bus stop is placed in a previously less connected neighborhood. In this scenario only one bus line is affected which limits the effect on other locations. Together with changes in the mobility profiles, our tool enables fast and accessible public transport planning.
@@ 4 Case Study
Our system is aimed primarily at policymakers and non-specialists who have an interest in the analysis and improvement of public transport infrastructure but do not have access to more detailed closed data and do not have the resources to work with numerical flow simulation planning tools. Such users might include political activists, lobbying groups, journalists, or interested citizen scientists. We argue that a broader representation of interested and supporting tools can strengthen the quality of public discourse and decision-making.
@@ 4.2 Case 2: Evaluating the Effect of Infrastructure Changes
(G2, G3) But how could such change look like, and how can one use our approach towards modeling and recommending potential changes to the transport infrastructure? For this case study, we chose the smaller city of Konstanz, Germany, with a population of around 85,000 inhabitants.
Alex is a student at the local university and is often frustrated that his home takes a long time to reach. He wonders if there are possible improvement opportunities in his neighborhood. Similar to the previous case, Alex performs a heatmap analysis, where he notices that bus lines stop on the very edge side of his neighborhood and hypothesizes that rerouting a bus route through its center might improve travel times for many houses. After identifying possible locations, he switches to the editing mode tab (Subsection 3.4). In the editing mode, the user can choose to place elements such as POIs, bike sharing, or new transport infrastructure (Figure 6). Alex wants to evaluate the effect of a changed bus stop placed at a strategic location. He first selects the bus stop icon in the menu bar. The bus line overlay is activated, and he can click on an existing bus line to move one stop from this bus line. After clicking on a bus line, he moves the stop from the edge to the center (Figure 6.2/3). Now, the simulation is re-computed in two stages, indicated by a loading icon: Within a region constrained by the neighboring bus stop locations, walking times to the closest location are re-computed to consider the change. Simultaneously, the travel times from the new bus stop to the relevant POIs are recomputed (i.e., a new row to the travel-time matrix is computed), as described above. Finally, for all relevant points, the travel time estimates are updated and are displayed on the travel time map (Figure 6.5). To evaluate the effect of the change, the user can select a difference visualization, i.e., a heatmap that displays the differences in travel time estimates with and without the modification (compare to Figure 7). Alex tries out multiple placements in order to identify the best possible location.
@@
4.3 Initial User Study
We conducted a initial user study with eleven participants recruited from a university environment, aged between 21 to 32 years: six computer science (CS) PhD students (PHD1-6), three CS master students (CS1-3), one finished CS master (CS4), and one history student (HS1). Of those, ten had prior knowledge of visual analytics, and seven had previously worked with GIS tools. The study was conducted in a university lab in a controlled environment, with a short preliminary questionnaire, a short introduction, a task-based assessment using think-aloud, followed by NASA TLX and User Experience Questionnaire (UEQ) as well as a semi-structured interview at the end, taking in total about one hour. The setup was shown on a 14.2-inch display with a resolution of 3024x1964, using the web-based prototype.
@@ 5 Discussion and Future Work
The case studies have shown the comparative ease with which a scenario-based analysis of existing public transport infrastructure can be generated, the effects of change can be analyzed, and the insights that can be drawn from it, which is validated by the initial user study. In the following, we want to highlight some key findings before discussing the limitations of our approach.
A difference map visualization showing additional types of analysis possible with the mobility profile editor: (a) the comparison of connectivity at night-time compared to day-time (white: no difference, magenta: large difference) (b) advantage of car usage over reliance on public transport (white: no difference, green: large difference).
@@ 5.1 Findings and Lessons Learned
We outline how our system matches the three goals of (G1) visualizing the quality of existing infrastructure, (G2) enabling detailed analysis for different demographic groups and usage scenarios, (G3) simulate the approximate effectiveness of modifications in the case studies (see Section 4 and initial user study (see Section 4.3). This is achieved via simple yet expressive visualizations (R1), a simulation based on open-source real-world data (R2), which is easily adaptable (R3).
@@ 5.2 Limitations and Future Work
The modifications of the transport network does not consider secondary effects in our existing simulation framework. Right now, for example, when adding a new stop or changing the schedule, it would make sense to perform changes to timetables also of other lines, adapting their schedules, too, to ensure smooth connections and reduce waiting times. Progressive schedule changes are currently not fully supported because they can quickly explode and would necessitate complex adaptation of timetables and routes, requiring an almost complete (and slow) re-simulation of the whole area, which is not compatible with the interactive approach. Our system could be extended to integrate other agent-based simulation frameworks like MatSim [40] or Sumo [21] or utilize more computationally extensive agent-based simulations.
@@
Our system is designed to be accessible and user-friendly, especially for non-expert users. While we expect some time required to become familiar with all functionalities, we have already evaluated the usability for non-expert users in a initial user study with eleven participants, indicating its usability for individuals or interest groups. The usability of the approach could be investigated in a quantitative user study. This could open up public transport planning to wider audiences and improve the quality of discussion for integrating civil society into planning processes. We can imagine our approach to enable even broader participation of citizens, e.g., we could use the mobility profile editor to collect survey data from citizens and aggregate the results to get more accurate profiles in areas where surveys have been limited so far.
The interactive nature requires some computing capability for a user, which is achievable for a domain expert setting but becomes a scalability issue when making the system open to a general audience. In the future, one could imagine a public read-only version of the system, while interactivity would be limited to a restricted user group, for example, through prior registration and application. Due to similar reasons, multi-modal and cross-modal transport is not yet considered, for example, combining public transport with individual mobility like a car. Our approach so far only considers public transport in itself, which might not always reflect actual usage (e.g., taking the bike to the train station), but this could be added with more compute available. When used by city planners and other expert users, integration with de-facto standard frameworks like ArcGIS might also be desirable.